OpenStudy (anonymous):

help!! #1,2a+3b-a,#2 (4xt3y)-(2x-5y),,#3,-3(2x-1)-5,,#4 -3(2x-5)-4(x+7)

OpenStudy (anonymous):

\[2a+3b-a=2a-a+3b=a+3b\]